if you have a string 13 foot long and its cut in two and the second half is 4 foot longer then the first half, how long is the two sections cut
Answer by macston(5194)

Let the shorter length=x, the longer=x+4
The total length (both pieces)=(x)+(x+4)=13 feet
2x+4=13 ft subtract 4 from each side
2x=9 ft divide each side by 2
x=4.5 ft ANSWER 1: The shorter piece is 4.5 feet.
longer=x+4=8.5 ft ANSWER 2: The longer piece is 8.5 feet.
CHECK
The total length=longer + shorter
13 feet=8.5 feet + 4.5 feet
13 feet=13 feet